OpenStudy (anonymous):

A blood bank needs twenty people to help with a blood drive. Twenty-five people have volunteered. Find how many different groups of twenty can be formed from the twenty-five volunteers.

OpenStudy (anonymous):

\[\dbinom{25}{20}\] or 25 choose 20 easier to compute the same number as \[\dbinom{25}{5}\] but the computation is still annoying

OpenStudy (anonymous):

\[\dbinom{25}{5}=\frac{25\times 24\times 23\times 22\times 21}{5\times 4\times 3\times 2}\] divide first, multply last, or use a calculator

OpenStudy (anonymous):

i get 53130 http://www.wolframalpha.com/input/?i=25+choose+20